Ok so i've done this problem at least 20 times and i don't know what i am doing wrong! The work seems to be right, but every time that i have done it, it does not check out properly. Please help!
Question: If the square of a positive number is decreased by twice the number the result is 15. Find the number.
Thank you!

Answer by stanbon(75887)

If the square of a positive number is decreased by twice the number the result is 15. Find the number.
--------------
Equation:
x^2-2x = 15
---
x^2-2x-15 = 0
Factor:
(x-5)(x+3) = 0
---
Positive solution:
x = 5
====================
Check:
5^2-2*5 = 15
25-10 = 15
15 = 15
====================
Cheers,
Stan H.
---